OpenStudy (amonoconnor):

Alright, this may seem extremely dumb, and elementary, but I would like some clarification as to the derivative of... cos^2(2x) 1. Can that be re-written as: (cos(2x))^2 ? Any and all help is greatly appreciated! :)

OpenStudy (mathstudent55):

Yes

OpenStudy (mathstudent55):

You can then separate the factors and use the derivative of a product. \(f(x) = \cos^2 2x = (\cos 2x)^2 = (\cos 2x)(\cos 2x) \)

OpenStudy (mathstudent55):

Do you know how to calculate the derivative of a product?

OpenStudy (amonoconnor):

Okay, then would the derivative be taken like such: d/dx (cos(2x))^2 = 2(cos(2x)) * (-sin(2x)) * 2

OpenStudy (mathstudent55):

\(f(x) = g(x)h(x)\) \(f'(x) = g(x)h'(x) + h(x) g'(x) \)

OpenStudy (amonoconnor):

Oh, does it need expanded like that to take the derivative?

OpenStudy (mathstudent55):

No. Anyway you can do it is fine.

OpenStudy (amonoconnor):

So is: d/dx (cos(2x))^2 = 2(cos(2x)) * (-sin(2x)) * 2 correct?

OpenStudy (mathstudent55):

\(2(\cos(2x)) * (-\sin(2x)) * 2\) You got the line above. Now you can use a trig identity to make it look simpler.

OpenStudy (amonoconnor):

Or... -4(cos(2x)^2)(sin(2x)

OpenStudy (mathstudent55):

Right, at least do this: \(-4\sin 2x \cos2x \)

OpenStudy (amonoconnor):

Is that the final derivative?

OpenStudy (mathstudent55):

Now you can do one more thing, if you want it to be even simpler looking. Use the identity \(\sin 2x = 2 \sin x \cos x\)

OpenStudy (amonoconnor):

Ohhh, so.... -2sin(4x) ?? :D

OpenStudy (mathstudent55):

\(-4\sin 2x \cos2x\) \(=-2 (2\sin 2x \cos2x) \) \(= -2 \sin 4x\)

OpenStudy (amonoconnor):

Is that accurate?

OpenStudy (mathstudent55):

Yes, you beat me to it. :-)

OpenStudy (amonoconnor):

Yay! You are an amazing helper... Thank you for actually working with me, not just regurgitating textbook terminology at this high-school brain.

OpenStudy (amonoconnor):

*thank you very much for the medal as well :)

OpenStudy (mathstudent55):

You are very welcome. I can tell you know your stuff. Good job!